OpenStudy (anonymous):

Find the product of (x + 7)(x - 7). x^2 - 14x + 49 x^2 - 49 x^2 + 14x + 49 x^2 + 49

OpenStudy (anonymous):

are you familiar with the FOIL method ?

OpenStudy (anonymous):

Foil stands for first, outside, inside, and last. (x + 7)(x - 7) First -- take the product of the first numbers in each set ... x * x = x^2 Outside -- product of outside numbers ... x * - 7 = -7x Inside -- product of inside numbers ... 7 * x = 7x Last -- product of last numbers ...7 * -7 = -49 put it all together... x^2 - 7x + 7x - 49 (the 7x's cancel each other out) leaving you with : x^2 - 49